Two resistors with resistors 20 ohms and 60 ohms are connected partly in series and partly in parallel. Draw the circuits with a DC power source and calculate the respective exchange resistance:

a) 20 ohm series with 60 ohms.

 b) 20 ohms in parallel with 60 ohms


1
Expert's answer
2022-05-02T13:41:54-0400
"R_a=R_1+R_2=20+60=80\:\Omega"

"R_b=\frac{R_1R_2}{R_1+R_2}=\frac{20*60}{20+60}=1.5\:\Omega"